What is the activity of a 0.05 mol sample of barium-137? (to 2 s.f and in Bq)
1
Expert's answer
2020-05-28T13:13:18-0400

Number of nuclei in  0.05 mol sample of barium-137 is

"N=0.05 mol(\frac{6.022*10^{23}}{1 mol})=0.3011*10^{23}"

Half life of barium-137 is "T_{1\/2}=153 s"

The activity of sample of barium-137 is

"R=\lambda N=(\frac{ln2}{T_{1/2}})N"

"=(\frac{ln2}{T_{1/2}})N=(\frac{ln2}{153 s})(0.3011*10^{23})=1.4*10^{20} Bq"